I need to factor to determine the zeros of the equation
Y=2xsquared +4x-16

Answer by algebrahouse.com(1659)

y = 2x² + 4x - 16
2x² + 4x - 16 = 0 {set y equal to zero, in order to find the zeros (x-intercepts)}
2(x² + 2x - 8) = 0 {factored 2 out of left side}
x² + 2x - 8 = 0 {divided each side by 2}
(x + 4)(x - 2) = 0 {factored left side into two binomials}
x + 4 = 0 or x - 2 = 0 {set each factor equal to 0}
x = -4 or x = 2 {solved each equation for x}

zeros are -4 and 2 as shown in the graph, below

For more help from me, visit: www.algebrahouse.com



.